Services Status: 3DOWN

IS&T Home > 3-Down

What you’ll find on this page:
General Services
Network
E-mail
Telephone Service
Academic Services
Library Services
Administrative Services

Greetings. Today is Saturday, January 21, 2017. This "3DOWN" Services Status page lists the status of the major services available to the MIT community, including scheduled and unscheduled outages.

We have had 5880 hits today with 5860 unique visitors.

General Services
Last Update: Monday, January 16, 2017 01:20 PM

Request Tracker (help.mit.edu) will be unavailable between 6pm and 7pm on Sunday, January 8th for routine maintenance.

January 14 2017 05:45AM: Planned SAP Production Outage on Saturday, January 14. The SAP production system was unavailable on Saturday, January 14 2017, between 01:30AM and 05:00AM. SAP, ECAT, and all Atlas SAP activity, including time sheet entry/approval, training, requisitioning, journal vouchers, credit card verification, events planning and service requests were inaccessible during this maintenance period. SAP production services are now available in accordance with their normal schedules.

Back To Top
Network
Last Update: Saturday, January 21, 2017 10:37 AM

MIT Legacy Cable TV system notice - WGBH signal on channels 2-1 and 2-2 is no longer operating. You can view the same content on channels 2-11 and 2-12. We do not have sufficient equipment in the legacy cable TV system to carry the new NBC affiliate station at this time.

Xfinity On Campus - The new NBC affiliate is now available on Channel 10 on the Xfinity On Campus service.

1/21/2017 10am: We've been made aware that there is an issue with fiber optic connections at MIT Lincoln Labs. This is also affecting connectivity to Haystack Observatory. Lincon network staff are working the issue. 

1/17/17 6:30-7:00: [COMPLETED] Building-68 will experience wired and wireless downtime due to scheduled Facilities power shutdown.

1/20/2017 6:00AM - 9:00AM: Building-46 wired and wireless network service will be down to service areas of 46-6018T due to Facilities normal/emergency power shutdowns.

1/20/17 6:30AM - 8:00AM EST:  [COMPLETED]  IS&T will be moving the APs in OC25 to a dedicated AP Group. A brief 5-10 min outage is expected for each AP in OC25 during the maintenance window.


 

Back To Top
Email
Last Update: Thursday, January 19, 2017 10:53 AM

January 19 2017 10:28AM: IS&T has resolved the issue with OWA.  We apologize for the inconvenience. 

Back To Top
Telephone Service
Last Update: Tuesday, January 3, 2017 10:45 AM

All services are operating normally.

Back To Top
Academic Services
Last Update: Thursday, January 5, 2017 07:59 AM

All services are operating normally.

Back To Top
Athena Message of the Day
Last Update: Friday, September 30, 2016 07:56 PM
On October 8, 2016 the 38-370 Athena cluster will be closed permanently. The Athena cluster combo will be changed on October 19, 2016. Please run: tellme combo at an Athena prompt to see the new combo.
Back To Top
Library Services
Last Update: Monday, September 26, 2016 12:57 PM

We are aware of a problem with the display of MIT-specific links in Google Scholar

ScienceDirect (Elsevier): Access appears to be restored. If you continue to experience problems, please contact us at Ask Us .
Back To Top
Administrative Services
Last Update: Saturday, January 14, 2017 05:44 AM

January 14 2017 05:45AM: Planned SAP Production Outage on Saturday, January 14. The SAP production system was unavailable on Saturday, January 14 2017, between 01:30AM and 05:00AM. SAP, ECAT, and all Atlas SAP activity, including time sheet entry/approval, training, requisitioning, journal vouchers, credit card verification, events planning and service requests were inaccessible during this maintenance period. SAP production services are now available in accordance with their normal schedules.

 

All services are operating normally.

Previous Versions

Follow the link below for the previous version of the Services Status page.
Saturday, January 21, 2017 10:37 AM